OpenStudy (anonymous):

A cylindrical water trough has a diameter of 6 ft and a height of 2 feet. It is being filled at the rate of 4 ft^3/min, so how fast is the water level rising when the water is 1 foot deep?

OpenStudy (anonymous):



jigglypuff314 (jigglypuff314):

V = pi r^2 h given info >,< V' = 4 h = 1 r' = 0 d = 6 so r = 3 looking for h' = ? anyways, first find the Derivative of V = pi r^2 h

OpenStudy (anonymous):

so, 0

jigglypuff314 (jigglypuff314):

wait... ... ... but dat not make sense >,< idk DX

OpenStudy (anonymous):

I'm confused. :(

jigglypuff314 (jigglypuff314):

I'm confused too sorry >,< @ranga can you help please???

jigglypuff314 (jigglypuff314):

@amoodarya could you help please?

jigglypuff314 (jigglypuff314):

@agent0smith @annas could you help please?

jigglypuff314 (jigglypuff314):

@shamil98 could you help please?

jigglypuff314 (jigglypuff314):

@agent0smith can you help? please?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

@austinL

OpenStudy (agent0smith):

Your problem is your drawing. It's oriented wrongly. That's not a trough. This is. Think of a horse or pig's trough 

OpenStudy (anonymous):

alright but how woud i go about solving the problem? the picture on my homework is a cylinder like a swimming pool

OpenStudy (agent0smith):

find dh/dt V = pi r^2 h you didn't differentiate w.r.t. time which is why you got zero.

OpenStudy (agent0smith):

\[\Large \frac{ dV }{ dt } = \pi r^2 \frac{ dh }{ dt }\]

jigglypuff314 (jigglypuff314):

why wouldn't it be V' = 2 pi r r' h' ?

OpenStudy (agent0smith):

r isn't changing (how can it, it's a constant radius throughout the vertical cylinder)

jigglypuff314 (jigglypuff314):

ohhhhhh right k

OpenStudy (anonymous):

jiggly, i don't get the one we're workin' on

OpenStudy (agent0smith):

\[\Large \frac{ dV }{ dt } = \pi r^2 \frac{ dh }{ dt }\]solve for dh/dt. You know dv/dt and r.

OpenStudy (agent0smith):

Guessing you solved this....?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

I still need help